Suppose that a professional wrestler is selected at random among 90 wrestlers, where 35 are over 350 pounds, 20 are bad guys, and 15 are over 350 pounds and bad guys. What is the probability that the wrestler selected is over 350 pounds or a bad guy?


1
Expert's answer
2021-11-11T18:32:22-0500

"P(over 350\cup bad)=P(over 350)+P(bad)-P(over 350\cap bad)="

"=\frac{35}{90}+\frac{20}{90}-\frac{15}{90}=\frac{40}{90}=\frac{4}{9}\approx0.4444."